# New Perspectives in Advancing Graph Machine Learning

### v1

@@ 2.3 Equivariant GNNs
Traditionally, both the message-passing networks and the geometric scattering transforms have mostly focused on abstract graphs, where the $v_{i}$ are merely thought of as elements of a set. However, for many applications, it is more natural to consider geometric graphs, where the vertices are points in some Euclidean space $\mathbb{R}^{d}$ and we are given both scalar-valued and vector-valued node features. For instance, with molecular graphs, the vertices may correspond to atoms with known coordinates in $\mathbb{R}^{3}$. Typically, one has access to both scalar-valued features, such as an encoding of the atom types, as well as vector-valued features such as atomic positions or velocities.
Formally, we define a geometric graph as a weighted, undirected graph $G=(V,E,w)$, where the vertices $v_{1},\dots,v_{n}$ lie in $\mathbb{R}^{d}$, and will denote the $i$-th coordinate of the vertex $v_{j}$ by $v_{j}[i]$. We will assume that we are given scalar-valued signals $\mathbf{x}_{i}:V\rightarrow\mathbb{R}$, $1\leq i\leq F_{\text{scalar}}$, and vector-valued signals $\mathbf{w}_{i}:V\rightarrow\mathbb{R}^{d}$, $1\leq i\leq F_{\text{vec}}$. Analogous to Section 2.1, we note that we may organize the scalar-valued signals into an $n\times F_{\text{scalar}}$ matrix $\mathbf{X}$; and similarly, we may organize the vector-valued features into an $n\times F_{\text{vec}}\times d$ tensor $\mathbf{W}$. However, for most of this paper, it will be convenient to work with individual signals $\mathbf{x}_{i}$ and $\mathbf{w}_{i}$.
When designing neural networks for geometric graphs, with both scalar-valued and vector-valued signals, it is important to preserve their intrinsic symmetries with respect to rotations. To understand this, consider what should happen if we rotate the entire system by some rotation matrix, $\mathbf{R}\in\mathbb{R}^{d\times d}$. Let $\mathbf{x}=\mathbf{x}^{(0)}$ denote an initial scalar-valued signal, and let $\mathbf{w}=\mathbf{w}^{(0)}$ denote an initial vector-valued singal. Let $\mathbf{x}^{(\ell)}$ and $\mathbf{w}^{(\ell)}$ denote our transformed representation after $\ell$ layers. Rotations do not affect scalar-valued signals such as atomic types. Thus, we aim to ensure that $\widetilde{\mathbf{x}}^{(\ell)}=\mathbf{x}^{(\ell)}$, where $\widetilde{\mathbf{x}}^{(\ell)}$ is analog of $\mathbf{x}^{(\ell)}$ but for the rotated system. This property is referred to as rotational invariance. However, for a vector-valued signals, such as the position or velocity of each vertex, rotating the system will also rotate the value of the feature. Therefore, we aim to ensure that $\widetilde{\mathbf{w}}^{(\ell)}=\mathbf{R}\mathbf{w}^{(\ell)}$. This property is referred to as rotational equivariance.
Equivariant GNNs (Satorras et al., 2021; Batzner et al., 2022; Han et al., 2022; Duval et al., 2023) are designed to possess these desirable properties. As an illustrative example, we consider EGNN (Satorras et al., 2021), and let $\mathbf{X}^{(\ell)}$ (and $\mathbf{W}^{(\ell)}$) be the matrix (tensor) of scalar-valued (vector-valued) features after the $\ell$-th layer. E-GNN first computes messages $m_{i,j}$ between adjacent nodes $v_{i}$ and $v_{j}$ that depend on $\mathbf{X}^{(\ell)}[i,:]$ and $\mathbf{X}^{(\ell)}[j,:]$ as well as the set of distances between the vector-valued features at $v_{i}$ and $v_{j}$, $\{\|\mathbf{W}^{(\ell)}[i,k,:]-\mathbf{W}^{(\ell)}[j,k,:]\|_{2}^{2}\}_{k=1}^{F_{\text{vec}}}$, which implies that $m_{i,j}$ is invariant under rotations. [^2] We present a slightly more general formulation than Satorras et al. (2021), which primarily focuses on the case when there is a single vector-valued signal representing the coordinates of the vertices in 33D space. It then (i) updates the scalar-valued node features to produce new features so that $\mathbf{X}^{(\ell+1)}[i,:]$ depends on the previous value $\mathbf{X}^{(\ell)}[i,:]$ as well as the messages $m_{i,j}$ and (ii) updates the vector-valued features to $\mathbf{W}^{(\ell+1)}[i,:,:]$ that depends on the previous value $\mathbf{W}^{(\ell)}[i,:,:]$ as well as "updates" of the form $\phi(m_{i,j})\left(\mathbf{W}^{(\ell)}[i,:,:]-\mathbf{W}^{(\ell)}[j,:,:]\right)$. Here $\phi$ is a scalar-valued function, which allows one to verify that these updates will commute with rotations.
Additionally, we note that unlike several previous works that introduce geometric scattering transforms (Chew et al., 2022, 2024; Johnson et al., 2025a) or more general GNNs (Wang et al., 2023, 2024; Chew et al., 2023) for geometric graphs, we do not necessarily assume that our graph is a discretization of an underlying manifold. Furthermore these works did not consider vector-valued node features.
3 Vector Diffusion Wavelets and Equivariant Scattering Transforms
@@
We construct a novel form of diffusion wavelets for vector-valued signals. This will allow us to define equivariant geometric scattering transforms for geometric graphs with both vector-valued and scalar-valued signals. As in Section 2.3, we will assume that we are given $F_{\text{scalar}}$ scalar-valued signals $\mathbf{x}_{i}:V\rightarrow\mathbb{R}$, $1\leq i\leq F_{\text{scalar}}$, and $F_{\text{vec}}$ vector-valued signals $\mathbf{w}_{i}:V\rightarrow\mathbb{R}^{d}$, $1\leq i\leq F_{\text{vec}}$. Importantly, we note that our framework is not restricted to the case that $d=3$, although this is the case for many real-world applications. As in Section 2.2, when convenient, we will identify signals with vectors. We will let $\mathbf{w}_{i}[j]=\mathbf{w}_{i}(v_{j})\in\mathbb{R}^{d}$ denote the value of the signal $\mathbf{w}_{i}$ at each point and we will view $\mathbf{w}_{i}$ as a vector in $\mathbb{R}^{nd}$ given by $\mathbf{w}_{i}=[\mathbf{w}_{i}[1]^{\top},\ldots,\mathbf{w}_{i}[n]^{T}]^{\top}=[\mathbf{w}_{i}[1][1],\ldots,\mathbf{w}_{i}[1][d],\ldots,\mathbf{w}_{i}[n][1],\ldots,\mathbf{w}_{i}[n][d]]^{\top}.$
In order to process the scalar-valued features $\mathbf{x}_{i}$, we may use the scalar-valued geometric scattering transform introduced in Section 2.2. However, in order to process the vector-valued features, we need to introduce a new form of geometric scattering. Towards this end, we will first define a vector diffusion matrix $\mathbf{Q}\in\mathbb{R}^{nd\times nd}$ using methods inspired by vector diffusion maps (Singer and Wu, 2012). We will then use $\mathbf{Q}$ to define vector diffusion wavelets and associated scattering networks.
@@
For each node $v_{i}\in\mathbb{R}^{d}$, we will construct a local basis $\{\mathbf{u}_{i,1},\ldots,\mathbf{u}_{i,d}\}$, which will provide a local coordinate system for each node relative to its neighbors. To build this local basis, we let $\mathcal{N}_{v_{i}}=\{v_{j}\in V:\{v_{i},v_{j}\}\in E\}$ denote the one-hop neighborhood of $v_{i}$ and let $n_{i}=|\mathcal{N}_{v_{i}}|$ denote the number of neighbors. [^3] We will assume that we have $n_{i}\geq d$ for all $i$. Otherwise, we will modify the edge set by adding edges between each $v_{i}$ and its nearest neighbors until $\text{deg}(v_{i})\geq d$. We then define a relative distance matrix $\mathbf{C}_{i}\in\mathbb{R}^{d\times n_{i}}$, whose columns are by $(v_{i_{j}}-v_{i})$ where $v_{i_{j}}$ is the $j$-th neighbor of $v_{i}$. In order to give more weight to nearer neighbors, we then rescale $\mathbf{C}_{i}$ by defining $\mathbf{B}_{i}=\mathbf{C}_{i}\mathbf{D}_{i},$ where $\mathbf{D}_{i}$ is a diagonal matrix defined by $\mathbf{D}_{i}[j,j]=\sqrt{K_{\epsilon}(v_{i},v_{i_{j}})}$, with $K_{\epsilon}(\cdot,\cdot)$ being a Guassian kernel with scale $\epsilon,$ $K(v_{i},v_{i_{j}})=\exp(-\|v_{i}-v_{i_{j}}\|_{2}^{2}/\epsilon)$.
@@
We next compute the singular value decomposition (SVD), $\mathbf{B}_{i}=\mathbf{U}_{i}\boldsymbol{\Sigma}_{i}\mathbf{V}_{i}^{\top}$. We note that, by the definition of the SVD, the matrix $\mathbf{U}_{i}$ is unitary, and so its columns $\{\mathbf{u}_{i,1},\dots,\mathbf{u}_{i,d}\}$ form an orthonormal basis for $\mathbb{R}^{d}$, which we interpret as a defining a local coordinate system centered around node $v_{i}$. We will assume throughout that each of the singular values, i.e., the diagonal entries of $\Sigma_{i}$, is in decreasing order and that none of the singular values have multiplicity greater than one.
@@
For all $i$ and $j$, we define $\mathcal{O}_{i,j}=\mathbf{U}_{i}\mathbf{U}_{j}^{\top}\in\mathbb{R}^{d\times d}$ which shifts between the local coordinate systems centered at $v_{i}$ and $v_{j}$. We note that the SVD is only unique up to sign-flips. That is, one may obtain a different SVD, $\mathbf{B}^{\prime}_{i}=\mathbf{U}^{\prime}_{i}\boldsymbol{\Sigma}_{i}(\mathbf{V}^{\prime})_{i}^{\top}$ by replacing both $\mathbf{u}_{i,k}$ and $\mathbf{v}_{i,k}$ with $\mathbf{u}^{\prime}_{i,k}=-\mathbf{u}_{i,k}$ and $\mathbf{v}^{\prime}_{i,k}=-\mathbf{v}_{i,k}$ for any fixed $k$. However, in order to ensure that the matrices $\mathcal{O}_{i,j}$ are well-defined (i.e., independent of these sign choices), we will employ a sign-flipping technique detailed in Appendix C.
@@
$\mathbf{Q}[i,j]=\mathbf{P}[i,j]\mathcal{O}_{i,j}\in\mathbb{R}^{d\times d}$
and observe that by construction, we have $\mathbf{Q}[i,j]=\mathbf{0}$ unless $\{v_{i},v_{j}\}\in E$. This implies we will only need to compute $\mathcal{O}_{i,j}$ if $\{v_{i},v_{j}\}\in E$ or $i=j$, which significantly reduces our computational cost.
Given $\mathbf{Q}$, we may then define vector diffusion wavelets and scattering coefficients analogous to (1) and (2). Specifically, we define $\widetilde{\mathcal{W}}_{J}=\{\widetilde{\mathbf{\Psi}}_{j}\}_{j=0}^{J}\cup\{\widetilde{\mathbf{\Phi}}_{J}\}$ by $\widetilde{\mathbf{\Psi}}_{0}=\mathbf{I}-\mathbf{Q}$, $\widetilde{\mathbf{\Phi}}_{J}=\mathbf{Q}^{2^{J}}$, and
$\widetilde{\mathbf{\Psi}}_{j}$ $=\mathbf{Q}^{2^{j-1}}-\mathbf{Q}^{2^{j}}=\mathbf{Q}^{2^{j-1}}(\mathbf{I}-\mathbf{Q}^{2^{j-1}}),\quad 1\leq j\leq J.$ (3)
@@
$\widetilde{\mathcal{U}}[j]\mathbf{w}(v)=\sigma(\widetilde{\boldsymbol{{\Psi}}}_{j}\mathbf{w}(v)),\quad\text{and}\quad\widetilde{\mathcal{U}}[j,j^{\prime}]\mathbf{w}(v)=\widetilde{\mathcal{U}}[j^{\prime}]\widetilde{\mathcal{U}}[j]\mathbf{w}(v)=\sigma(\widetilde{\boldsymbol{{\Psi}}}_{j^{\prime}}\sigma(\widetilde{\boldsymbol{{\Psi}}}_{j}\mathbf{w}(v))).$ (4)
As in the scalar-feature case, if one wishes, $m$-th order scattering coefficients for $m\geq 3$ can be defined similarly the rule $\widetilde{\mathcal{U}}[j_{1},j_{2},\ldots,j_{m}]\mathbf{w}(v)=\widetilde{\mathcal{U}}[j_{m}]\widetilde{\mathcal{U}}[j_{1},\ldots,j_{m-1}]\mathbf{w}(v).$ Additionally, when it is convenient to think of the signal $\mathbf{w}$ as a vector in $\mathbb{R}^{nd}$, we will write the first-order coefficients as $\widetilde{\mathcal{U}}[j]\mathbf{w}=\sigma(\widetilde{\boldsymbol{{\Psi}}}_{j}\mathbf{w})$ instead of $\widetilde{\mathcal{U}}[j]\mathbf{w}(v)=\sigma(\widetilde{\boldsymbol{{\Psi}}}_{j}\mathbf{w}(v))$ (and similarly with the higher-order coefficients).
@@ 3.1 Theoretical Results
The following result establishes frame bounds for the vector diffusion wavelets, similar to analogous results for the scalar-valued diffusion wavelets (Gama et al., 2018; Perlmutter et al., 2023). It implies that vector diffusion wavelet transform may be stably inverted and is robust to additive noise. [^4] Proofs of all theorems are provided in Appendix A.
@@
$c\>\frac{d_{\min}}{d_{\max}}\|\mathbf{w}\|_{2}^{2}\leq\big\|\widetilde{\mathcal{W}}_{J}\mathbf{w}\big\|^{2}_{2}\coloneqq\sum_{j=0}^{J}\big\|\widetilde{\boldsymbol{\Psi}}_{j}\mathbf{w}\big\|^{2}_{2}+\big\|\widetilde{\boldsymbol{\Phi}}_{J}\mathbf{w}\big\|^{2}_{2}\leq\frac{d_{\max}}{d_{\min}}\|\mathbf{w}\|^{2}_{2}.$
We next show that our vector diffusion wavelets and scattering coefficients are equivariant to the actions of the Special Orthogonal group, $SO(d)$, i.e., the set of all $d\times d$ rotation matrices. Will assume that our entire system has been subjected to the same global rotation $\mathbf{R}\in SO(d)$ and use bars to denote objects in the rotated coordinate system so that e.g., $\overline{v}_{i}=(\overline{v}_{i}[1],\overline{v}_{i}[2],\overline{v}_{i}[3])^{\top}$ denotes the position of the $i$-th vertex in the rotated system and $\overline{\mathbf{Q}}$ denotes the vector diffusion matrix constructed from the $\overline{v}_{i}$. We observe that we have $\overline{v}_{i}=\mathbf{R}v_{i}$, and, motivated by examples such as when our vector-valued-node features are either the coordinates or the velocities of each vertex, we will assume that rotating the entire system rotates the values of vector-valued features so that we have $\overline{\mathbf{w}}_{f}=\mathbf{R}\cdot\mathbf{w}$, $\mathbf{R}\cdot\mathbf{w}_{f}=[(\mathbf{R}\mathbf{w}_{f}(v_{1})^{\top},\ldots,(\mathbf{R}\mathbf{w}_{f}(v_{n}))^{\top}]^{\top}.$ Furthermore, we will assume that rotations do not change the connectivity structure of the graph so that we have $\overline{\mathbf{A}}=\mathbf{A}$. The following theorems establish the equivariance of vector diffusion wavelets and associated scattering coefficients. We also note Figure 2 in the appendix, which provides a visual illustration of Theorem 3.2.
@@
For any vector-valued node feature $\mathbf{w}$, we have, a
$\overline{\widetilde{\mathbf{\Psi}}}_{j}\overline{\mathbf{w}}=\overline{\widetilde{\mathbf{\Psi}}}_{j}(\mathbf{R}\cdot\mathbf{w})=\mathbf{R}\cdot\widetilde{\mathbf{\Psi}}_{j}\mathbf{w},\quad\text{for all}\quad 0\leq j\leq J,\quad\text{and}\quad\overline{\widetilde{\mathbf{\Phi}}}_{J}\overline{\mathbf{w}}=\mathbf{R}\cdot\widetilde{\mathbf{\Phi}}_{j}\mathbf{w}.$
@@
$\overline{\widetilde{\mathcal{U}}}[j_{1},\ldots,j_{m}]\overline{\mathbf{w}}=\overline{\widetilde{\mathcal{U}}}[j_{1},j_{2},\ldots,j_{m}](\mathbf{R}\cdot\mathbf{w})=\mathbf{R}\cdot\widetilde{\mathcal{U}}[j_{1},\ldots,j_{m}]\mathbf{w}.$
We note that Theorem 3.3 introduces the assumption that $\sigma$ commutes with rotations. To understand why this condition is necessary, consider the case where $m=1$ and recall the definition, $\widetilde{\mathcal{U}}[j_{1}]\mathbf{w}(v)=\sigma(\widetilde{\boldsymbol{\Psi}}_{j_{1}}\mathbf{w}(v))$. Applying Theorem 3.2 allows us to see that
Therefore, the equivariance result will not hold without this assumption. We also note that we may readily construct activations $\sigma$ which satisfy this assumption by defining $\sigma(\mathbf{w})(v)=\sigma_{\text{radial}}(\|\mathbf{w}(v)\|_{2})\frac{\mathbf{w}(v)}{\|\mathbf{w}(v)\|_{2}}$ when $\mathbf{w}(v)\neq 0$, and $\sigma(\mathbf{w})(v)=0$ when $\mathbf{w}(v)=0$, where $\sigma_{\text{radial}}$ is any real-valued function defined on $0,\infty)$.

### v2

@@ 2.3 Equivariant GNNs
In this section, we consider geometric graphs where the vertices $\{v_{i}\}_{i=1}^{n}\subseteq\mathbb{R}^{D}$ are not necessarily assumed to lie upon a $d$-dimensional manifold, $d<D$. We note that even in this case, we can still design a network which aims to utilize intrinsic structure and symmetries of the data.
Taking inspiration from Equivariant GNNs (Satorras et al., 2021; Batzner et al., 2022; Han et al., 2022; Duval et al., 2023), in Section 3, we will show that our network has desirable symmetries with respect to rotations. Specifically, we let $\mathbf{R}\in\mathbb{R}^{D\times D}$ denote a rotation matrix that we assume acts on our vertex set $\{v_{i}\}_{i=1}^{n}\rightarrow\{\mathbf{R}v_{i}\}_{i=1}^{n}$. We let $\mathbf{x}=\mathbf{x}^{(0)}$ and $\mathbf{w}=\mathbf{w}^{(0)}$ denote initial scalar-valued and vector-valued signals, and we let $\mathbf{x}^{(\ell)}$ and $\mathbf{w}^{(\ell)}$ denote our representations after $\ell$ layers.
The manner in which our network processes the scalar-valued signals will not be affected by the action of the rotation $\mathbf{R}$, i.e., we have that $\overline{\mathbf{x}}^{(\ell)}=\mathbf{x}^{(\ell)}$, where $\overline{\mathbf{x}}^{(\ell)}$ is analog of $\mathbf{x}^{(\ell)}$ but for the rotated system. This property is referred to as rotational invariance. For the vector-valued signals, we will show that rotating the system rotates our represtentation in the corresponding manner, i.e., $\overline{\mathbf{w}}^{(\ell)}=\mathbf{R}\mathbf{w}^{(\ell)}$. This property is referred to as rotational equivariance. Equivariant GNNs such as (Satorras et al., 2021; Batzner et al., 2022; Han et al., 2022; Duval et al., 2023) have shown that these symmetries act as an effective inductive bias which facilitates the learning process.
3 Vector Diffusion Wavelets and VDW-GNNs
@@
As in Section 2, we assume that we are given $F_{\text{scalar}}$ scalar-valued signals $\mathbf{x}_{i}:V\rightarrow\mathbb{R}$, $1\leq i\leq F_{\text{scalar}}$, and $F_{\text{vec}}$ vector-valued signals $\mathbf{w}_{i}:V\rightarrow\mathbb{R}^{D}$, $1\leq i\leq F_{\text{vec}}$. As in Section 2.1, when convenient, we will identify signals with vectors. We let $\mathbf{w}_{i}[j]=\mathbf{w}_{i}(v_{j})\in\mathbb{R}^{D}$ denote the value of the signal $\mathbf{w}_{i}$ at each vertex, and we when view $\mathbf{w}_{i}$ as a vector in $\mathbb{R}^{nD}$ we write $\mathbf{w}_{i}=[\mathbf{w}_{i}[1]^{\top},\ldots,\mathbf{w}_{i}[n]^{T}]^{\top}=[\mathbf{w}_{i}[1][1],\ldots,\mathbf{w}_{i}[1][D],\ldots,\mathbf{w}_{i}[n][1],\ldots,\mathbf{w}_{i}[n][D]]^{\top}.$
@@
To process the vector-valued features, we introduce a new form of diffusion wavelets, which are inspired by vector diffusion maps (Singer and Wu, 2012) and defined in term of a vector diffusion matrix $\mathbf{Q}\in\mathbb{R}^{nD\times nD}$. For each node $v_{i}\in\mathbb{R}^{D}$, we construct a local basis $\{\mathbf{u}_{i,1},\ldots,\mathbf{u}_{i,D}\}$, which provides a local coordinate system for each node relative to its neighbors. To build this local basis, we let $\mathcal{N}_{v_{i}}=\{v_{j}\in V:\{v_{i},v_{j}\}\in E\}$ denote the one-hop neighborhood of $v_{i}$ and let $n_{i}=|\mathcal{N}_{v_{i}}|$ denote the number of neighbors. [^2] We assume that we have $n_{i}\geq D$. Otherwise, we add edges between each $v_{i}$ and its nearest neighbors until $\text{deg}(v_{i})\geq D$. We then define a relative distance matrix $\mathbf{C}_{i}\in\mathbb{R}^{D\times n_{i}}$, whose columns are denoted by $(v_{i_{j}}-v_{i})$ where $v_{i_{j}}$ is the $j$-th neighbor of $v_{i}$. In order to give more weight to nearer neighbors, we then rescale $\mathbf{C}_{i}$ by defining $\mathbf{B}_{i}=\mathbf{C}_{i}\mathbf{D}_{i},$ where $\mathbf{D}_{i}$ is a diagonal matrix defined by $\mathbf{D}_{i}[j,j]=\sqrt{K_{\epsilon}(v_{i},v_{i_{j}})}$, with $K_{\epsilon}(\cdot,\cdot)$ being a Guassian kernel with scale $\epsilon,$ $K(v_{i},v_{i_{j}})=\exp(-\|v_{i}-v_{i_{j}}\|_{2}^{2}/\epsilon)$.
@@
We next compute the singular value decomposition (SVD), $\mathbf{B}_{i}=\mathbf{U}_{i}\boldsymbol{\Sigma}_{i}\mathbf{V}_{i}^{\top}$. We note that, by the definition of the SVD, the matrix $\mathbf{U}_{i}$ is unitary, and so its columns $\{\mathbf{u}_{i,1},\dots,\mathbf{u}_{i,D}\}$ form an orthonormal basis for $\mathbb{R}^{D}$, which we interpret as a defining a local coordinate system centered around node $v_{i}$. We assume throughout that each of the singular values, i.e., the diagonal entries of $\Sigma_{i}$, are in decreasing order and that none of the singular values have multiplicity greater than one. This ensures that the SVD is unique up to sign flips of the singular vectors.
@@
For all $i$ and $j$, we define $\mathcal{O}_{i,j}=\mathbf{U}_{i}\mathbf{U}_{j}^{\top}\in\mathbb{R}^{D\times D}$ which shifts between the local coordinate systems centered at $v_{i}$ and $v_{j}$ as shown in Figure 1. We note that since the SVD is only unique up to sign-flips, one may obtain a different SVD, $\mathbf{B}^{\prime}_{i}=\mathbf{U}^{\prime}_{i}\boldsymbol{\Sigma}_{i}(\mathbf{V}^{\prime})_{i}^{\top}$ by replacing both $\mathbf{u}_{i,k}$ and $\mathbf{v}_{i,k}$ with $\mathbf{u}^{\prime}_{i,k}=-\mathbf{u}_{i,k}$ and $\mathbf{v}^{\prime}_{i,k}=-\mathbf{v}_{i,k}$ for any fixed $k$. However, in order to ensure that the matrices $\mathcal{O}_{i,j}$ are well-defined (i.e., independent of these sign choices), we employ a sign-flipping technique that, when needed, replaces $\mathbf{u}_{i,k}$ with $\mathbf{u}_{i,k}$ in order to ensure that $\langle\mathbf{u}_{i,k},\mathbf{u}_{j,\ell}\rangle$ is always non-negative (full details are provided in Appendix C).
Figure 1: An example point cloud of data lying on a continuous manifold. Dashed lines: edges of a $k$-NN graph. Solid green lines: neighbor edges used to generate the local PCA for each node. Blue arrows: local coordinate systems/frames $U_{i}$ and $U_{j}$ induced by the local graph structure. $\mathcal{O}_{i,j}$ is the rotation that transforms $U_{j}$ to $U_{i}$.
@@
$\mathbf{Q}[i,j]=\mathbf{P}[i,j]\mathcal{O}_{i,j}\in\mathbb{R}^{D\times D}.$
We observe that for $i\neq j$, we have $\mathbf{Q}[i,j]=\mathbf{0}$ unless $\{v_{i},v_{j}\}\in E$ in which case we do not need to compute $\mathcal{O}_{i,j}$. This significantly reduces our computational cost. (Details on computational complexity are given in Appendix D.)
Given $\mathbf{Q}$, we may then define vector diffusion wavelets and scattering coefficients analogous to (2), (3), and (4). Specifically, we define $\widetilde{\mathcal{W}}_{J}=\{\widetilde{\mathbf{\Psi}}_{j}\}_{j=0}^{J}\cup\{\widetilde{\mathbf{\Phi}}_{J}\}$ by
$\widetilde{\mathbf{\Psi}}_{j}$ $=\mathbf{Q}^{2^{j-1}}-\mathbf{Q}^{2^{j}}=\mathbf{Q}^{2^{j-1}}(\mathbf{I}-\mathbf{Q}^{2^{j-1}}),$ (5)
for $1\leq j\leq J$, $\widetilde{\mathbf{\Psi}}_{0}=\mathbf{I}-\mathbf{Q}$, and $\widetilde{\mathbf{\Phi}}_{J}=\mathbf{Q}^{2^{J}}$.
@@
$\widetilde{\mathcal{U}}[j]\mathbf{w}(v)=\sigma(\widetilde{\boldsymbol{{\Psi}}}_{j}\mathbf{w}(v)),\quad\text{and}$ (6)
$\widetilde{\mathcal{U}}[j,j^{\prime}]\mathbf{w}(v)\hskip-0.72229pt=\hskip-0.72229pt\widetilde{\mathcal{U}}[j^{\prime}]\widetilde{\mathcal{U}}[j]\mathbf{w}(v)\hskip-0.72229pt=\hskip-0.72229pt\hskip-0.72229pt\sigma(\widetilde{\boldsymbol{{\Psi}}}_{j^{\prime}}\sigma(\widetilde{\boldsymbol{{\Psi}}}_{j}\mathbf{w}(v))).$ (7)
When it is convenient to think of the signal $\mathbf{w}$ as a vector in $\mathbb{R}^{nd}$, we will write the first-order coefficients as $\widetilde{\mathcal{U}}[j]\mathbf{w}=\sigma(\widetilde{\boldsymbol{{\Psi}}}_{j}\mathbf{w})$ instead of $\widetilde{\mathcal{U}}[j]\mathbf{w}(v)=\sigma(\widetilde{\boldsymbol{{\Psi}}}_{j}\mathbf{w}(v))$ (and similarly with the higher-order coefficients).
Analogous to, e.g., Min et al. (2020; 2021); Tong et al. (2024); Johnson et al. (2025a); Wenkel et al. (2024); Viswanath et al. (2025), in our experiments we will use these vector-diffusion wavelets and scattering transforms as the basis for variety of GNNs, which we refer to as Vector Diffusion Wavelet Graph Neural Networks (VDW-GNNs). Additionally, we note in the scalar-valued signal case, Tong et al. (2024) showed that the scattering transform could be extended to include non-dyadic diffusion wavelets of the form $\mathbf{P}^{t_{j}}-\mathbf{P}^{t_{j+1}}$ where $0=t_{0}<t_{1}<t_{2}\ldots$ is a generic sequence of increasing integers and that vector diffusion wavelets may be readily extended to non-dyadic scales in exactly the same way.
@@ 3.1 Theoretical Results
The following result establishes frame bounds for the vector diffusion wavelets, similar to results for the scalar-valued diffusion wavelets (Gama et al., 2018; Perlmutter et al., 2023). It implies that the vector diffusion wavelet transform may be stably inverted and is robust to additive noise. [^3] Proofs of all theorems are provided in Appendix A.
@@
$c\>\frac{d_{\min}}{d_{\max}}\|\mathbf{w}\|_{2}^{2}\leq\big\|\widetilde{\mathcal{W}}_{J}\mathbf{w}\big\|^{2}_{2}\leq\frac{d_{\max}}{d_{\min}}\|\mathbf{w}\|^{2}_{2},$
where $d_{\min}$ and $d_{\max}$ denote the minimal and maximal vertex degrees and $\big\|\widetilde{\mathcal{W}}_{J}\mathbf{w}\big\|^{2}_{2}\coloneqq\sum_{j=0}^{J}\big\|\widetilde{\boldsymbol{\Psi}}_{j}\mathbf{w}\big\|^{2}_{2}+\big\|\widetilde{\boldsymbol{\Phi}}_{J}\mathbf{w}\big\|^{2}_{2}$.
We next show that our vector diffusion wavelets and scattering coefficients are equivariant to the actions of the Special Orthogonal group, $SO(D)$, i.e., the set of all $D\times D$ rotation matrices. We assume that our entire system has been subjected to the same global rotation $\mathbf{R}\in SO(D)$ and use bars to denote objects in the rotated coordinate system so that e.g., $\overline{v}_{i}=(\overline{v}_{i}[1],\overline{v}_{i}[2],\ldots,\overline{v}_{i}[D])^{\top}$ denotes the position of the $i$-th vertex in the rotated system and $\overline{\mathbf{Q}}$ denotes the vector diffusion matrix constructed from the $\overline{v}_{i}$. We observe that we have $\overline{v}_{i}=\mathbf{R}v_{i}$, and, motivated by examples such as when our vector-valued-node features are, e.g., the coordinates of each vertex, we assume that rotating the entire system rotates the values of vector-valued features so that we have $\overline{\mathbf{w}}_{f}=\mathbf{R}\cdot\mathbf{w}$, $\mathbf{R}\cdot\mathbf{w}_{f}=[(\mathbf{R}\mathbf{w}_{f}(v_{1})^{\top},\ldots,(\mathbf{R}\mathbf{w}_{f}(v_{n}))^{\top}]^{\top}.$ Furthermore, we assume that rotations do not change the connectivity structure of the graph so that we have $\overline{\mathbf{A}}=\mathbf{A}$. The following theorems establish the equivariance of vector diffusion wavelets and associated scattering coefficients. We also note Figure 4 in the appendix, which provides a visual illustration of Theorem 3.2.
@@
For any vector-valued node feature $\mathbf{w}$, we have, $\overline{\widetilde{\mathbf{\Phi}}}_{J}\overline{\mathbf{w}}=\mathbf{R}\cdot\widetilde{\mathbf{\Phi}}_{j}\mathbf{w}$ and
$\overline{\widetilde{\mathbf{\Psi}}}_{j}\overline{\mathbf{w}}=\overline{\widetilde{\mathbf{\Psi}}}_{j}(\mathbf{R}\cdot\mathbf{w})=\mathbf{R}\cdot\widetilde{\mathbf{\Psi}}_{j}\mathbf{w},\quad\text{for all}\quad 0\leq j\leq J.$
@@
$\overline{\widetilde{\mathcal{U}}}[j_{1},\ldots,j_{m}]\overline{\mathbf{w}}$ $=\overline{\widetilde{\mathcal{U}}}[j_{1},j_{2},\ldots,j_{m}](\mathbf{R}\cdot\mathbf{w})$
$=\mathbf{R}\cdot\widetilde{\mathcal{U}}[j_{1},\ldots,j_{m}]\mathbf{w}.$
We note that Theorem 3.3 introduces the assumption that $\sigma$ commutes with rotations, which we discuss in Appendix A.3. Additionally, we note that since our relative distance matrices $\mathbf{C}_{i}$ defined in terms of the $v_{i_{j}}-v_{i}$, our entire method is automatically completely translation invariant.